OpenStudy (anonymous):

a rectangular bathtub with a base 6 by 4 feet is filled with water to a height of 1 foot. a dog is placed in the water for a bath and the water rises 1/2 foot. what is the dogs volume?

jimthompson5910 (jim_thompson5910):

Volume the water in the tub: V = LWH V = 6*4*1 V = 24 cubic feet -------------------------- Since the water rose 1/2 foot, this means that the new height of the water is now 1+1/2 = 2/2 + 1/2 = 3/2 feet So the new volume of the water is V = LWH V = 6*4*(3/2) V = 24*3/2 V = 72/2 V = 36 Now subtract the two volumes: New Volume - Old Volume 36 - 24 12 So the dog's volume is 12 cubic feet

OpenStudy (anonymous):

thank you i need help with another question

jimthompson5910 (jim_thompson5910):

you're welcome, whats the question?